ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 62, 92 Marrero, Osvaldo, Villanova University, USA Statistical analysis of seasonality in multiple groups This work is concerned largely with the statistical analysis of seasonal variation in epidemiological data, which usually have small sample size or small amplitude, or both. Against such data, standard statistical procedures often have inadequate power. Two statistical procedures are presented and discussed. One is a test for seasonality that can detect simultaneously different kinds of seasonal variation in different groups. The other is a test for homogeneity of seasonal patterns of the same kind among different groups. The procedures are illustrated with examples.